Fermi National Laboratory


Accelerator Update for August 26, 2005 - August 29, 2005

Friday, August 26
The day shift began with Operations monitoring store 4349, the Recycler Ring (RR) stash at 141.5 E10, the Antiproton Source (Pbar) stack at 152.2 E10, and with MiniBooNE, NuMI, and the Meson MCenter taking beam.

Store 4349 was terminated at 10:41 AM.

Shot setup begins at 12:19 PM.

Operations established store 4351 at 2:54 PM with an initial luminosity of 113.22 E30.

Main Injector (MI) began Pbar acceleration studies at 3:39 PM.

At 5:22 PM Pbar acceleration studies are complete. Pbar stacking resumes, beam returns to MCenter.

Operations put the Recycler into shot setup at 7:52 PM to stash antiprotons.

Pbar returns to stacking at 8:59 PM, 10 minutes later beam returns to MCenter.

Saturday, August 27
The midnight shift began with Operations monitoring store 4351, the Recycler Ring (RR) stash at 74.7 E10, the Pbar stack at 54.5 E10, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations put the Recycler into shot setup at 12:40 AM to stash antiprotons.

Pbar returns to stacking at 1:17 AM, beam to MCenter.

Store 4351 was terminated at 1:05 PM.

Both CDF and D0 went into controlled access at 1:20 PM.

At 2:05 PM problems with Linac quads Q51-Q54 kept beam off everywhere while experts diagnosed and repaired the problem. This lasted until 7:08 PM.

Shot setup begins at 7:22 PM.

Operations established store 4353 at 9:53 PM with an initial luminosity of 106.38 E30.

Pbar returns to stacking.

Sunday, August 28
The midnight shift began with Operations monitoring store 4353, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations put the Recycler into shot setup at 12:34 AM to stash antiprotons.

Pbar returns to stacking at 1:15 AM.

At 1:19 AM we received a power glitch when feeder 46A tripped off. When Feeder 46A tripped it took out various Cryo compressors around the Tevatron ring, also lost some network connections, which dropped the QPM link, thus resulting in the loss of store 4353. The glitch also caused us to lose beam to other areas.

Beam returns to NuMI, MiniBoone, and MCenter at 2:22 AM. We also returned to stacking at this time.

At 5:42 AM feeder 46A was re-energized, in order to do this the air breaker at A3 had to be bypassed. The air breaker was damaged due to a flashover caused by a mouse being in the wrong place at the wrong time. The Cryo compressors are now back on, but will take some time to recover (~ 6 hours).

Operations put the Recycler into shot setup at 6:07 AM to stash antiprotons.

Pbar returns to stacking at 6:37 AM, beam returns to MCenter.

Tevatron turn on began at 1:00 PM.

The Tev performed a dry squeeze at 3:04 PM, followed by a wet squeeze at 3:38 PM.

Shot setup begins at 4:29 PM.

At 5:15 PM we received another power glitch when feeder 47 tripped off. The glitch caused many problems in Linac and Booster, it also caused a few problems in the Tevatron. Shot setup has been delayed until we recover from the glitch.

At 7:47 PM feeder 47 was re-energized after isolating from the cause of the trip (road and grounds and Site 37).

At 8:52 PM, Tevatron recovery from the glitch is complete, and the Tev is in the first of 3 HEP ramps.

After Linac and Booster problems were resolved, beam returns to NuMI, MiniBoone, and MCenter at 11:47 PM.

Monday, August 29
The midnight shift began with Operations preparing for shot setup.

Shot setup begins at 1:07AM.

Operations established store 4358 at 2:40 AM with an initial luminosity of 111.74 E30.

Operations put the Recycler into shot setup at 6:06 AM to stash antiprotons.

Pbar returns to stacking at 6:54 AM, beam returns to MCenter.

The Plan for Monday
Keep the store, continue stacking and shooting to the recycler, run beam to NuMI, MiniBoone, and MCenter.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/29/2005   email Fermilab